OpenStudy (anonymous):

what is the value of c so that x^2 - 11x + c is a perfect square trinomial?

OpenStudy (anonymous):

a) 121 b) 121/4 c) -11/2 d) 121/2

OpenStudy (anonymous):

what is half of eleven?

OpenStudy (anonymous):

as a fraction

OpenStudy (anonymous):

its supposed to be in fraction form

OpenStudy (anonymous):

yes, i know

OpenStudy (anonymous):

yeah sorry for the confusion

OpenStudy (anonymous):

so what is half of eleven?

OpenStudy (anonymous):

ohhh you're asking me

OpenStudy (anonymous):

yes, i am asking what half of eleven is as a fraction

OpenStudy (anonymous):

5.5

OpenStudy (anonymous):

as a fraction

OpenStudy (anonymous):

not a decimal or a mixed number, just a fraction

OpenStudy (anonymous):

i dont know to be honest

OpenStudy (anonymous):

then i will tell you half of eleven is \(\frac{11}{2}\)

OpenStudy (anonymous):

to get your answer, square that one. i.e. compute \[\frac{11}{2}\times \frac{11}{2}\]

OpenStudy (anonymous):

so 121/4 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohh thanks!

OpenStudy (anonymous):

yw it always works that way